# Evaluating language models as risk scores

### v1

@@ A.3 Varying uncertainty
for an in-depth description of each categorical value a feature can take.
As expected, predictive signal increases when using larger feature sets, but not in a straightforward way: some features actually worsen performance. This contrasts with standard supervised learning algorithms, where all features would either be neutral or contribute positively to a model’s predictive power when tested on data that is i.i.d. with the training data. On the other hand, there is no clear trend for calibration: for Mistral models, it seems that calibration actually worsens for larger feature sets, while for Llama models calibration is approximately stable across all points.
@@ A.4 Feature importance
The folktexts package is made available to the public via its open-source code repository1 and as a standalone package to be installed via the Python Package Index (PyPI). The main user-facing classes are Benchmark, BenchmarkConfig, LLMClassifier, TaskMetadata, ColumnToText, and Dataset. The responsibilities of each class are ascribed as follows
@@
usage: run_acs_benchmark [-h] model MODEL task-name TASK_NAME results-dir RESULTS_DIR data-dir DATA_DIR [–few-shot FEW_SHOT] [–batch-size BATCH_SIZE] [–context-size CONTEXT_SIZE] [–fit-threshold FIT_THRESHOLD] [–subsampling SUBSAMPLING] [–seed SEED] [–dont-correct-order-bias] [–chat-prompt] [–direct-risk-prompting] [–reuse-few-shot-examples] [–use-feature-subset [USE_FEATURE_SUBSET …]]
[–use-population-filter [USE_POPULATION_FILTER …]] [–logger-level {DEBUG,INFO,WARNING,ERROR,CRITICAL}]
Run a benchmark folktexts task.
@@
task-name TASK_NAME
[str] Name of the ACS task to run the experiment on
@@
chat-prompt [bool] Whether to use chat-based prompting (for instruct models)
direct-risk-prompting
[bool] Whether to directly prompt for risk-estimates instead of multiple-choice Q&A
@@
use-feature-subset [USE_FEATURE_SUBSET …]
[str] Optional subset of features to use for prediction
use-population-filter [USE_POPULATION_FILTER …]
@@
Figure A6: Documentation for using folktexts package through the command-line interface. An executable named run_acs_benchmark is made available to run the standard ACS benchmark tasks.

### v3

@@ A.4 Varying uncertainty
for an in-depth description of each categorical value a feature can take. Predictive signal (ROC AUC) reliably increases with each added feature, for all tested models except the Gemma 2B variants. This is expected with standard supervised learning algorithms trained and evaluated on i.i.d. data, but arguably somewhat unexpected of pre-trained models trained on a variety of datasets that are out-of-distribution relative to the evaluation set. On the other hand, there is no clear trend for calibration: for Mistral models, it seems that calibration actually worsens for larger feature sets, while for Llama models calibration is approximately stable across all points.
@@ A.5 Feature importance
The folktexts package is made available to the public via its open-source code repository1 and as a standalone package to be installed via the Python Package Index (PyPI). It is compatible with PyTorch models used locally, as well as with web-hosted models available through an API. The main user-facing classes are Benchmark, BenchmarkConfig, LLMClassifier, TaskMetadata, ColumnToText, and Dataset. The responsibilities of each class are ascribed as follows
@@
usage:
run_acs_benchmark [-h] model MODEL results-dir RESULTS_DIR data-dir DATA_DIR [–task TASK] [–few-shot FEW_SHOT] [–batch-size BATCH_SIZE] [–context-size CONTEXT_SIZE] [–fit-threshold FIT_THRESHOLD] [–subsampling SUBSAMPLING] [–seed SEED] [–use-web-api-model] [–dont-correct-order-bias] [–numeric-risk-prompting] [–reuse-few-shot-examples] [–use-feature-subset USE_FEATURE_SUBSET]
[–use-population-filter USE_POPULATION_FILTER] [–logger-level {DEBUG,INFO,WARNING,ERROR,CRITICAL}]
Benchmark risk scores produced by a language model on ACS data.
@@
use-web-api-model [bool] Whether use a model hosted on a web API (instead of a local model)
@@
numeric-risk-prompting
[bool] Whether to prompt for numeric risk-estimates instead of multiple-choice Q&A
@@
use-feature-subset USE_FEATURE_SUBSET
[str] Optional subset of features to use for prediction, comma separated
use-population-filter USE_POPULATION_FILTER
@@
Figure A10: Documentation for using folktexts package through the command-line interface. An executable named run_acs_benchmark is made available to run the standard ACS benchmark tasks with a variety of available customization options. Detailed documentation available at socialfoundations.github.io/folktexts/